# Activation Steering for Robust Type Prediction in CodeLLMs

### v1

@@ 4.1 Layer Ablations
We evaluate the accuracy of steering on each edit dataset (table 1). To find the optimal layer for applying steering vectors, we conduct an ablation on TypeScript data with StarCoderBase-1B. We consider patching at single layers as well as intervals of three and five adjacent layers. Figure 4 shows that for code edits where steering is most effective (e.g., renaming types), patching at single versus multiple layers makes little difference. However, for the less effective edits (e.g., removing type annotations), patching multiple layers significantly increases accuracy. We hypothesize that this performance gap exists because task vector refinement occurs over multiple layers, thus needs to be steered over multiple layers. This follows from previous ideas that transformers build predictions incrementally throughout layers (Elhage et al., 2021; Geva et al., 2022). Following this finding, we chose to patch layers 10–14 in StarCoderBase-1B and 19–23 in StarCoderBase-7B.
@@
Table 2: Steering StarCoderBase-1B and 7B on several datasets of semantics-preserving code edits. For each dataset, we report the accuracy on a held out set of negative prompts, i.e., the model mispredicts all of these types without steering and steering corrects a significant number of mispredictions. The Random column reports accuracy after steering with a randomly initialized steering vector.
4.2 Evaluating Different Edits and Effectiveness By Type
Figure 5: Mean accuracy of steering Python by expected type label frequency. Error bars indicate the interquartile range. Note that we constrain the maximum size of any class to 25.
Having picked the layers to steer, table 2 summarizes the results of steering on each of our models. For StarCoderBase-1B, steering is most effective at fixing distractors from renaming types, achieving an accuracy of 90% in Python and 75% in TypeScript. We report accuracy over the held-out evaluation sets for each edit type. However, we also evaluated steering accuracy on the steering set and found no significant difference in performance between steering splits or held-out splits (appendix B). This is a first indication that our steering vectors generalize well.
@@ 4.2 Evaluating Different Edits and Effectiveness By Type
4.3 Backup Circuits
A threat to validity for any intervention that involves activation patching is that the patch may not be directly improving performance, but just triggering a backup circuit (McGrath et al., 2023). For example, it could be that patching is just introducing noise into the embedding space that causes other mechanisms to trigger, thus producing the expected outcome. This effect complicates the interpretability of patches as well as steering vectors.
As part of our evaluation, we steer the model with a random steering vector (the Random columns in table 2). We find that steering in this way has non-zero accuracy, though it is generally significantly lower than our computed steering vectors. We posit that the non-zero accuracy of the random vector occurs due to backup circuits. However, the more significant gains from our computed steering vectors indicate that our steering method performs true transformations towards the correct type.
4.4 Steering Versus Finetuning
Fine-tuning is the traditional approach to improving model performance on a downstream task, so we now compare steering to fine-tuning. We fine-tune StarCoderBase-1B on the Python Renamed Types dataset. We fine-tune for five epochs using the AdamW optimizer and a learning rate of $3\times 10^{-5}$ with weight decay of $0.1$. At the end of each epoch, we evaluate on the held-out set.
In figure 6(a), we see that fine-tuning is approximately as effective as steering on the type prediction task. However, figure 6(b) shows that this comes with a caveat: fine-tuning on type-prediction significantly degrades the model’s code completion ability, as measured with HumanEval Chen et al. (2021). In contrast, since the steering vectors for type prediction are a small, localized patch, one can easily toggle them on or off based on the task at hand. On the other hand, fine-tuning for type prediction produces a new specialized model. We have shown that this fine-tuned model is weak at code completion, but in general, other abilities may have been impacted too.
(a) Type prediction accuracy.
(b) HumanEval pass@1.
Figure 6: We fine-tune StarCoderBase-1B on the Python Renamed Types dataset. We find that within 1 epoch, the fine-tuned model achieves the same accuracy on the held-out set as the steered model (figure 6(a)). However, this fine-tuning makes the model much worse at code completion (figure 6(b)).
4.5 Language Transfer in Steering Vectors
@@
Syntactically, Python and TypeScript look very different. However, semantically the two languages have a lot in common Politz et al. (2013); Bierman et al. (2014). In particular, both of them use gradual typing disciplines with many shared semantic features. So, we wondered, could CodeLLMs be learning a representation of types that is shared between these two languages?
To test this hypothesis, we test if steering vectors built on TypeScript data can improve the accuracy of Python type prediction, and vice versa. We conduct this experiment with each of our datasets: we steer StarCoderBase-1B using vectors from language $A$ but evaluate on the corresponding held-out test set from language $B$. Table 3 shows that this is nearly as effective as steering type prediction on the same language.
From this observation, we hypothesize that within the model there exists a representation of program types that is shared across TypeScript and Python. This is in contrast to the scenario where distinct model components are specialized to each programming language. Thus, steering vectors are effective across languages because they intervene on the shared representation between them. Moreover, these results suggest that steering vectors generalize by targeting a high-level representation of types.
@@ 5 Conclusion
We investigate activation steering for type prediction by making CodeLLMs more robust against semantically-irrelevant aspects of code. We find that by constructing steering pairs using semantics-preserving code edits we can construct highly effective steering vectors. Our experiments show that steering vectors generalize outside the programs used for steering and outperform naive baselines. We find that our steering vectors are transferable across languages. This suggests the existence of a shared representation of a type within the CodeLLM.
Activation steering can be a powerful technique for improving model performance on tasks where fine-tuning is not advantageous. As CodeLLMs are trained on more programming languages and downstream tasks, activation steering may be used as a lightweight alternative to multiple fine-tuned experts. Rather than fine-tuning several models on different tasks, the same base model could be specialized through steering vectors which are added and removed as needed. This could be especially convenient for resource-constrained applications. Our type prediction vectors, for example, could make a convenient lightweight expert for a coding assistant, useful for applications like type migration.
In future work, we aim to study the underlying causal mechanism in CodeLLMs responsible for type prediction. We further wish to explore how type-prediction steering vectors may generalize to open generation problems like code completion.
6 Ethical Considerations
The purpose of this work is to investigate methods for making CodeLLMs more robust and aligned to user intent. It is our view that interpreting CodeLLMs is necessary for understanding whether models approach programming in a principled way. As models become more integrated into developers’ workflows, model errors could compromise the security of entire systems. For this reason, we focus on making models more robust against distractors.
The datasets used in our investigation are sourced from publicly available code. Our TypeScript dataset is derived from a subset of The Stack v1.2, which contains permissively licensed data with personal identifying information (PII) filtered. The ManyTypes4Py dataset is funded by the European Commission, which follows data privacy laws under the EU General Data Protection Regulation (GDPR). We utilize the StarCoder family of models for their open and publicly available weights and training data.
7 Reproducibility
We commit to making all of our code and datasets public for the purpose of reproducibility and furthering research in the mechanistic interpretability of CodeLLMs.
@@ Acknowledgments
We thank Ming-Ho Yee for providing the TypeScript dataset that we use in this work. We thank Northeastern Research Computing for providing the computing resources used for this work. This work is partially supported by the U.S. National Science Foundation (CCF-2052696).

### v2

@@ 4.1 At What Layers is Steering Most Effective?
To find the optimal layer for activation steering, we conduct a layer ablation using TypeScript data with StarCoderBase-1B (figure 4). We evaluate the accuracy of steering on each edit dataset (table 1), and we explore patching at single layers as well as intervals of three and five adjacent layers.
There are two takeaways from this experiment. First, we find that steering middle and later layers is most effective. In the early layers of the model, the hidden state of the last token represents the last token itself. As computation moves through layers and layer outputs are incrementally added to the residual stream, this hidden state is refined into the next token prediction. This phase change in computation coincides with the model identifying the inference task, thus makes the target of our steering.
Secondly, we show that steering at multiple layers improves performance. Specifically, for code edits where steering is most effective (e.g., renaming types), patching at single versus multiple layers makes little difference. However, for the less effective edits (e.g., removing type annotations), patching multiple layers significantly increases accuracy. We hypothesize that this performance gap exists because task vector refinement occurs over multiple layers. In these cases, activations need to be steered over multiple layers. This follows from previous ideas that transformers build predictions incrementally throughout layers (Elhage et al., 2021; Geva et al., 2022).
@@
Table 2: Steering StarCoderBase-1B and 7B on datasets of semantics-preserving edits. For each dataset, we report the accuracy on a held out set of negative prompts, i.e., the model mispredicts all of these types before steering. The Random column reports accuracy after steering with a randomly initialized steering vector.
4.2 Does Steering Work Across Types and Different Kinds of Edits?
We now consider the effectiveness of steering for different kinds of edits. Based on the results of the previous section, we do so with patches applied to layers 10–14 for StarCoderBase-1B and 19–23 for StarCoderBase-7B.
Table 2 summarizes the results under the Steering columns. We report accuracy over held-out evaluation sets for each edit type. The effectiveness of steering can vary significantly. For StarCoderBase-1B, steering is most effective at fixing distractors from renaming types, achieving an accuracy of 90% for Python and 75% for TypeScript. (Note that without steering, the models mispredict on all problems in the dataset.) We also evaluate steering accuracy on the steering set and find no significant difference in performance between steering splits or held-out splits (appendix B). This is a first indication that our steering vectors generalize well.
@@ 4.2 Does Steering Work Across Types and Different Kinds of Edits?
Figure 5: Mean accuracy of steering Python by expected type label frequency. Error bars indicate the interquartile range. Each class has maximum size of 25.
4.3 How Does Steering Compare to Baselines?
Backup Circuits
A threat to validity for any intervention that involves activation patching is that the patch may not be directly improving performance, but just triggering a backup circuit (McGrath et al., 2023). For example, it could be that patching is just introducing noise into the embedding space that causes other mechanisms to trigger, thus producing the expected outcome. This effect complicates the interpretability of steering vectors.
For a baseline, we steer the model with a random steering vector (the Random columns in table 2). We find that steering this way has non-zero accuracy, but it is significantly lower than our computed steering vectors. Thus we conclude that the steering vectors we compute perform true, localized transformations towards the correct type prediction task in activation space.
@@ 4.3 How Does Steering Compare to Baselines?
Fine-tuning
It is not yet clear if the accuracy of our steering is “good”. To answer this question, we compare steering to fine-tuning. Intuitively, fine-tuning a model on a narrow task gives us an upper bound on how good the model can be at that task. We fine-tune StarCoderBase-1B on the Python Renamed Types dataset. We fine-tune for five epochs using the AdamW optimizer and a learning rate of $3\times 10^{-5}$ with weight decay of $0.1$. At the end of each epoch, we evaluate on the held-out set. We see that fine-tuning is approximately as effective as steering (figure 6). This further indicates that steering targets and steers the part of the model responsible for the task.
Figure 6: We fine-tune StarCoderBase-1B on the Python Renamed Types dataset. We find that within 1 epoch, the fine-tuned model achieves the same accuracy on the held-out set as the steered model.
4.4 Do Models Learn Type Prediction Across Languages?
Python and TypeScript are syntactically distinct, but their semantics have a lot in common Politz et al. (2013); Bierman et al. (2014). Both languages use gradual typing disciplines with many shared semantic features. So, could CodeLLMs be learning a mechanism for type prediction that is language-agnostic?
To test this hypothesis, we evaluate if steering vectors built on TypeScript data can improve the accuracy of Python type prediction, and vice versa. We conduct this experiment with each of our datasets: we steer StarCoderBase-1B using vectors from language $A$ but evaluate on the corresponding held-out test set from language $B$. Table 3 shows that this is nearly as effective as steering type prediction on the same language.
Collectively, results indicate that our steering vectors do not represent a specific type token like str or int. Instead, they encode a delta which steers model generations toward a mechanism for type prediction that 1) generalizes across different source codes; 2) is less sensitive to semantically irrelevant distractors; 3) generalizes across the languages we study. Given these observations, we conclude that there exists a non-spurious mechanism for type prediction in CodeLLMs which, when activated either natively or through activation steering, is more robust against distractors.
@@ 5 Conclusion
We investigate how CodeLLMs perform gradual type prediction using the two most widely used gradually typed languages: TypeScript and Python. We leverage activation steering and FIM prompting as interpretability tools to investigate what happens when a model mispredicts. By using semantics-preserving code edits, we construct distractors that cause the model to mispredict. We then show that we can recover performance by constructing steering vectors to guide the model towards more robust type prediction.
Our steering vectors are as effective as fine-tuning a model on type prediction. This indicates that our vectors effectively steer the part of the model responsible for type prediction. Furthermore, Python steering vectors are as effective at steering TypeScript, and vice versa. This provides new evidence that type prediction in CodeLLMs is not spurious, but a causal mechanism learned across languages that generalizes beyond syntax.
6 Limitations
Our findings shed light on how CodeLLMs perform the type prediction task for TypeScript and Python. Both these languages are well represented in CodeLLM training corpora. However, our findings may not generalize to low-resource gradually typed languages, e.g., Typed Racket Tobin-Hochstadt and Felleisen (2008) or Luau Lily Brown et al. (2023) since the performance of base models on these languages is very poor.
We show that our approach works for two sizes of the StarCoderBase model family. We chose these models because we found them to be more effective at type prediction than other models trained with FIM Since StarCoder training data is fully open, we know the quantity and quality of StarCoder FIM training is high–a necessary requirement for fairly evaluating steering on FIM type prediction.
7 Ethics Statement
The purpose of this work is to understand whether CodeLLMs perform type prediction using reliable mechanisms. It is our view that interpreting CodeLLMs is necessary for understanding whether models approach programming in a principled way. As CodeLLMs become more integrated into developers’ workflows, model errors could compromise the security of entire systems. For this reason, we focus on making models more robust against distractors.
The StarCoderBase models have an OpenRAIL license which allows the kind of investigation in this paper. Our TypeScript dataset is derived from a subset of The Stack v1.2, which contains permissively licensed data with personal identifying information (PII) filtered. The ManyTypes4Py dataset is funded by the European Commission, which follows data privacy laws under the EU General Data Protection Regulation (GDPR). These datasets are intended for CodeLLMs, which this paper investigates.
We utilize the StarCoder family of models for their open and publicly available weights and training data. The size and efficient implementation of attention in these models allows us to use limited compute resources. We were able to run experiments on a single A100-80GB GPU. In total, we estimate around 48h of GPU time, including preliminary experiments and fine-tuning. Note that the computation of our final steering vectors is a small fraction of this time.
@@ Acknowledgments
Portions of this work are implemented with NNSight and NDIF Fiotto-Kaufman et al. (2024). We thank Ming-Ho Yee for help with the TypeScript dataset that we use in this work Yee (2024). We thank Northeastern Research Computing for providing the computing resources used for this work.